# Charged excitations made neutral: N-centered ensemble density functional theory of Fukui functions

### v1

@@ Abstract
This is an astounding simplification: the only piece needed to compute Fukui functions aside from a ground-state functional without any approximation is the weight-derivatives the ensemble Hxc potential evaluated for null weights. In practice, one does not need to build an ensemble DFA covering the whole range of valid weights to obtain accurate results, but merely to model the weight dependence of the potential in the immediate vicinity of the familiar $N$-electron ground-state limit.
To that aim, it is appealing to ”dress” a regular KS-DFA with a weight-dependent scaling function
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s(\bm{\xi})\;v^{\bm{\xi}=0}_{\rm Hxc}[n]=s(\bm{\xi})\;v_{\rm Hxc}[n]$ (14)
where $s(\bm{\xi}=0)=1$ by construction. The weight-derivatives would then be approximated as
$\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|^{\bm{\xi}=0}_{n=n_{0}}\approx\frac{\partial s(\bm{\xi})}{\partial\xi_{\lambda}}\Big|_{\bm{\xi}=0}v_{\rm Hxc}[n_{0}]$ (15)
This takes full advantage of the 0-weight limit mindset, $s(\bm{\xi})$ only accounting for the ensemble behavior in the vicinity of $\bm{\xi}=0$, and benefits from all the known exact constraints built in the ground-state functional provided the scaling function is chosen in a way that complies with them. In particular, this choice complies with exact uniform coordinate scaling constraint valid both for regular KS DFA and for ensembles Nagy (1995); Scott et al. (2024). Known limits of the ensemble Hxc functional can be used to build $s(\bm{\xi})$, as we will explore in the following. Throughout this work, we will use the exact 2-electrons Hubbard ground-state Hxc functional in pair with scaling functions.
Low correlation regime – Orbital-based approximations were used extensively in EDFT to bypass development of explicit (weight-dependent) density functionals, the simplest being an Hartree-Exchange approximation to the ensemble energyYang et al. (2014); Pribram-Jones et al. (2014); Yang et al. (2017); Gould et al. (2018); Sagredo and Burke (2018); Deur et al. (2018); Senjean and Fromager (2018); Deur and Fromager (2019); Senjean and Fromager (2020). When the correlation is not too strong, one can expect such an approach to do well, which we put to the test by considering the exact exchange ensemble functional (EEXX) of the Hubbard Dimer. We recall it is obtained as the first order term in the Taylor expansion in $U$ of the universal functional,
$E^{\bm{\xi}}_{\rm Hx}[n]=U\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}$ (16)
and yields all quantities in Eq. (9) by differentiation of $E^{\bm{\xi}}_{\rm Hx}[n]$. All results are computed in the 0-weight limit.
Figure 1: Fukui functions in weak correlation regime for null ensemble weights (parameters in the red inset). Full and dashed lines are $f_{+}$ and $f_{-}$ calculations respectively. Exact results (yellow lines) are compared to EEXX (blue lines) and to the neglect of weight-derivative contributions (grey lines).
In the weak correlation regime $U/t=0.75$, the EEXX functional closely reproduces exact results as seen on Fig. 1, with the effect of neglecting Hxc weight derivatives shown in grey. The improvement is clear, making this first application of EEXX to compute densities promising.
These encouraging results suggest using the scaling function
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hxc}[n]\,,\qquad s_{\rm Hx}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}{v_{\rm Hx}[n^{\bm{\xi}}]}$ (17)
Note the ensemble density is fed to the ground-state potential. It is important to define the scaling in this way to ensure evaluating the weight-derivative of the scaling function (in the 0-weight limit) does not require knowing the weight-derivative of the ensemble density. While using this scaling function misses correlation at the ensemble level, note some measure of it is still brought into the approximate potential through the ground-state Hxc potential (see Eq. (17)). For $U/t=0.75$, there is negligible difference between the EEXX scaling method (not shown) and full EEXX DFA, which could be expected as correlation barely contributes in this regime.
Figure 2: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to EEXX scaling (red) and neglect of weight derivative terms (grey). Exact results are shown in orange.
Intermediate correlation regime – We expect this to change by turning to higher (but moderate) correlation $U/t=1.5$, results being shown on Fig. 2. For external potentials below $\Delta v=2U/t$, where the correlation energy becomes important, we see EEXX deviates from the exact result. It still improves substantially over the neglect of weight-derivative terms for $f_{-}$, but not for $f_{+}$: the weight-derivative terms only bring a small correction here, which is overestimated in EEXX. Turning to the scaling strategy, we ask: Is ground-state correlation a reasonable approximation of its ensemble analog? Results show dressing the ground-state correlation potential with a purely Hx ensemble scaling brings a modest improvement over full EEXX for $f_{-}$, but prediction of $f_{+}$ actually worsens.
These results show it is necessary to go beyond EEXX when correlation takes over. Moreover, we have to understand if the failure of Hx scaling function comes from an inherent discrepancy between ground-state and ensemble correlation, making the weight-dependent dressing strategy impractical, or if it can be fixed. To deal with situations where EEXX fails, second order (Görling-LeviGörling and Levy (1994)) perturbation theory has been applied to ensemble DFT, yielding an orbital-dependent DFA beyond EEXXYang (2021). We transpose this in the Hubbard dimer through second-order Taylor expansion in $U$:
$E^{\bm{\xi}}_{\rm c}[n]\overset{\rm PT2}{\approx}\frac{U^{2}}{2}\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}$ (18)
Compared to EEXX-only on Fig. 3, it shows much better agreement with exact Fukui functions.
Figure 3: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to PT2 (green) and neglect of weight derivative terms (grey). Exact results are shown in orange.
Moreover, comparison of functional forms of EEXX and PT2 shows where our Hx-only scaling is failing: The ratio of PT2 ensemble and ground-state correlation potentials is an explicit function of the weights and of the ensemble density:
$s_{\rm c,PT2}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm c,PT2}[n^{\bm{\xi}}]}{v_{\rm c,PT2}[n^{\bm{\xi}}]}=s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})$ (19)
while in the Hx ratio Eq. (17) all dependence on the ensemble density simplifies out, leaving only explicit weight dependences. This disparity in complexity for the Hx and correlation part suggests using a different scaling function for each, which reads at the PT2 level:
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hx}[n]+s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})\;v_{\rm c}[n]$ (20)
with exact ensemble density dependence shown explicitly. Note if the correlation term was neglected all-together, leaving only the first term, the potential would be equal to that of EEXX (in the 0-weight limit). But at the PT2 level, it dresses the exact ground-state correlation with an ensemble-density dependence. The question now becomes: Does this double scaling at the PT2 level improve over the full PT2 EDFA? On Fig. 4, we compare their performance in computing $f_{+}$ for $U/t$ up to 2.5. We can see that the PT2 scaling accurately matches exact results even above $U=1.5$, when both EEXX and PT2 fail. This shows the scaling function strategy holds promise, but it is crucial to treat the Hx and correlation parts separately.
Figure 4: Fukui function $f_{+}$ in the low correlation regime for null ensemble weights (parameters in the red inset). For increasing $U$, we compare our EEXX EDFA (blue) to PT2 EDFA (green) and the PT2 scaling strategy (red). Exact results (orange) and neglect of weight-derivatives (grey) are plotted for reference.
Strong correlation – Learning from high- and low-correlation limits of the EDFT proved to be an outstanding approach for excitation energies in difficult situationsGould et al. (2023), with a versatile EDFA obtained from interpolating between the two regimes. We put it to the test for Fukui functions by building a Padé interpolation between the PT2 (low $U$) and $U/t\rightarrow\infty$ limits of the ensemble Hxc functional,
$E_{\rm Hxc}^{\bm{\xi}}[n]=a^{\bm{\xi}}[n]\,U+\frac{b^{\bm{\xi}}[n]\,U^{2}}{1+c^{\bm{\xi}}[n]\,U}$ (21)

### v2

@@ Abstract
$f_{s,+}$ being the lowest unoccupied KS molecular orbital (LUMO) in the $N$-electron ground state, $f_{s,-}$ the HOMO, and $n_{0}:=n^{\bm{\xi}=0}=n_{\Psi_{0}}$. This is an astounding simplification where the only piece needed (aside from the regular $N$-electron Hxc kernel) to compute Fukui functions, in principle exactly, is the weight-derivatives of the Nc ensemble density-functional Hxc potential evaluated at $\bm{\xi}\equiv\{\xi_{+},\xi_{-}\}=0$. In practice, one does not need to build an EDFA covering the whole range of positive weight values to obtain accurate results, but merely to model the weight dependence of the Hxc potential in the immediate vicinity of the familiar $N$-electron ground-state limit. To that aim, it is appealing to “dress” any regular DFA with a weight-dependent scaling function,
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s(\bm{\xi})\;v^{\bm{\xi}=0}_{\rm Hxc}[n]=s(\bm{\xi})\;v_{\rm Hxc}[n],$ (14)
where $s(\bm{\xi}=0)=1$, by construction. In principle, one could use Hxc energies ratio to define $s(\bm{\xi})$, but as illustrated in the supplemental material, their weight-dependence differ widely from that of the potentials, making them unsuited for the task. Unlike a bare (weight-independent) DFA, the above EDFA will enable the description of weight-derivatives, as follows,
$\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|^{\bm{\xi}=0}_{n=n_{0}}\approx\frac{\partial s(\bm{\xi})}{\partial\xi_{\pm}}\Big|_{\bm{\xi}=0}v_{\rm Hxc}[n_{0}].$ (15)
While being useful, it finds its theoretical justification in the uniform coordinate scaling exact constraints that both DFA and EDFA are expected to fulfill Nagy (1995); Scott et al. (2024). Known limits of the exact ensemble Hxc functional can be used to build $s(\bm{\xi})$, for example in the weakly and strictly correlated regimes Gould et al. (2023). Considering the former, a sensible pick would consist in combining approximations based on the ensemble exact exchange (EEXX), which has been used extensively to compute excitation energies Yang et al. (2014); Pribram-Jones et al. (2014); Yang et al. (2017); Gould et al. (2018); Sagredo and Burke (2018); Deur et al. (2018); Senjean and Fromager (2018); Deur and Fromager (2019); Senjean and Fromager (2020), with a perturbative treatment of electron correlation at the ensemble level, as proposed in Ref. 59. A priori, this strategy cannot tackle strongly correlated systems. While the scaling approach of Eq. (14) may improve on this thanks to the partial account of correlation through the ground-state functional, a more global strategy that interpolates between extreme limits is needed to ensure that EDFA can describe all correlation regimes within the ensemble. This proved to be an outstanding approach for excitation energies in difficult situations Gould et al. (2023), for example. We show in the following how these different strategies can be implemented successfully for the computation of Fukui functions.
Illustrative example – As a proof of principle, we apply the present Nc EDFT formalism to the simple but nontrivial asymmetric Hubbard dimer Carrascal et al. (2015, 2016); Fromager (2020). The model allows to study all regimes of electronic correlation, by tuning the on-site (atomic) electronic repulsion to the electronic hopping (which is analogous to the kinetic energy strength) ratio $U/t$, for a given difference $\Delta v$ in external potential between the two sites, while the ensemble Hxc functional and its derivatives can be computed exactly. We focus on charged excitations between different electronic ground states, as described in Eq. (13), with the central number of electrons set to $N=2$. Exact ensemble densities are systematically used so that we can focus on functional-driven (and, in particular, weight-driven) errors. We start with the following EEXX-scaled functional, $v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})v_{\rm Hxc}[n]$, where the scaling function $s_{\rm Hx}(\bm{\xi})={v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}/{v_{\rm Hx}[n^{\bm{\xi}}]}$ relies on the (analytically solvable in the model) EEXX scheme. Note that, in the targeted zero-weight limit, the weight derivative of the EEXX scaling function does not involve the response of the ensemble density to weight variations, by construction. While reducing to regular EEXX when electron correlation is negligible, the above EDFA can in principle capture some correlations through the ground-state correlation potential contribution $s_{\rm Hx}(\bm{\xi})v_{\rm c}[n]$.
Figure 1: Performance of the EEXX-scaled EDFA in evaluating Fukui functions (as functions of the external asymmetry potential $\Delta v$) from the zero-weight limit of Nc EDFT in the moderately correlated $U/t=1.5$ regime of the Hubbard dimer.
Results are shown in Fig. 1 in the moderately correlated $U/t=1.5$ regime. While EEXX deviates from the exact result, as expected, it still improves substantially over the complete neglect of weight derivatives in the evaluation of the ionization Fukui function $f_{-}$, but not for the affinity one ($f_{+}$). In the latter case, ensemble correlation effects mostly compensate the EEXX ones. Unfortunately, the EEXX-scaled functional worsens the prediction of $f_{+}$, while bringing a modest improvement over EEXX for $f_{-}$. Therefore, it seems essential to correlate the scaling function. For that purpose, the extension to ensembles Yang (2021) of Görling–Levy perturbation theory Görling and Levy (1993, 1994) through second order (PT2) can be used. As shown in the supplementary material, unlike in EEXX, the density-functional correlation scaling function $v^{\bm{\xi}}_{\rm c,PT2}[n]/v_{\rm c,PT2}[n]=:s_{\rm c,PT2}[\bm{\xi},n]$ is both weight- and density-dependent in the model, thus suggesting the following refinement of the ansatz introduced in Eq. (14),
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hx}[n]+s_{\rm c,PT2}[\bm{\xi},n^{\bm{\xi}}]\;v_{\rm c}[n],$ (16)
with a separate scaling for Hx and correlation potentials. While PT2 (applied on top of EEXX) fixes the flaw of EEXX mentioned previously, when $U/t=1.5$, the PT2-scaled ansatz of Eq. (16) has the advantage of extending the applicability of PT2 (through the scaling function) to stronger correlation regimes, with a remarkably high accuracy, as illustrated in Fig. 2 for the affinity Fukui function. A deeper analysis of the correlation potential, which illustrates how the scaling approach corrects erroneous behaviors of PT2 for strongly asymmetric dimers, is provided in the supplemental material. It clearly shows that the scaling function strategy holds promises, but it is crucial to treat Hx and correlation parts separately.
Figure 2: Performance of the PT2-scaled EDFA in evaluating the affinity Fukui function (as a function of the interaction strength $U$) up to the stronger correlation $U/t=2.5$ regime of the asymmetric ($\Delta v/t=3.0$) Hubbard dimer.
For completeness, we also extended to charged excitations the ”interpolation between extreme limits” strategy of Ref. 37. In the Hubbard dimer, it translates into the following Padé approximant (see the supplementary material for the complete expression) to the ensemble density-functional Hxc energy,
$E_{\rm Hxc}^{\bm{\xi}}[n]\approx a^{\bm{\xi}}[n]{U}+\frac{b^{\bm{\xi}}[n]U^{2}}{1+\frac{b^{\bm{\xi}}[n]}{\gamma^{\bm{\xi}}[n]-a^{\bm{\xi}}[n]}U},$ (17)
which is exact in both weakly ($U/t\rightarrow 0$) and strictly correlated ($U/t\rightarrow+\infty$) limits, by construction, and from which the Fukui functions can be evaluated through density and weight derivatives. To be applicable (and numerically stable) to intermediate correlation regimes, a smoothed version of the coefficient $\gamma^{\bm{\xi}}[n]$ in Eq. (17) has been designed such that the Hxc kernel is reproduced at all interaction strengths in the pure ground-state ($\bm{\xi}=0$) and symmetric ($\Delta v=0$) case. The latter is the analog of the uniform electron gas in this context, echoing recent work to develop (semi-) local density functionals for excited states Loos (2025).
Figure 3: Same as Fig. 1 for the Padé approximant-based EDFA in the strongly correlated $U/t=5$ regime and away from the zero-weight limit ($\xi_{+}=\xi_{-}=0.2$).
The robustness of the smoothed Padé approximant away from the zero-weight limit it learned from (we take $\xi_{+}=\xi_{-}=\xi=0.2$) is illustrated in Fig. 3, where it successfully applies to the strongly correlated $U/t=5$ dimer in all asymmetry regimes, unlike EEXX, as expected.
In summary, an alternative ($N$-centered) EDFT of Fukui functions has been derived, thus revealing that important but challenging higher-order Hxc derivative discontinuities can be tackled, in principle exactly, through ensemble weight derivative corrections to regular DFT. As a result, significant progress made recently in the understanding and application of EDFT Gould et al. (2026, 2025) can now be exploited for computing accurately charged excited-state properties beyond the energy, as illustrated in this Letter.
This work has benefited from support provided by the University of Strasbourg Institute for Advanced Study (USIAS) for a Fellowship, within the French national programme “Investment for the future” (IdEx-Unistra).
Supplemental Material
Appendix S-I Analytical expressions for the Hubbard dimer
S-I.1 Exact expressions
Model – We consider a model diatomic system built from the Hubbard dimer, with Hamiltonian
$\hat{H}_{e}=-t\sum_{\sigma}\big(\hat{a}_{0\sigma}^{\dagger}\hat{a}_{1\sigma}+\hat{a}_{1\sigma}^{\dagger}\hat{a}_{0\sigma}\big)+U\sum_{i=0}^{1}\hat{n}_{i\uparrow}\hat{n}_{i\downarrow}+\frac{\Delta v}{2}\big(\hat{n}_{1}-\hat{n}_{0}\big)$ (S1)
[…]
$H_{N}=\begin{bmatrix}U-\Delta v&-\sqrt{2}\,t&0\\ -\sqrt{2}\,t&0&-\sqrt{2}\,t\\ 0&-\sqrt{2}\,t&U+\Delta v\end{bmatrix}$ (S2)
the exact energies being analytically given by the third-order polynomial equationCarrascal et al. (2015, 2016); Deur et al. (2017):
$E_{N,i}^{3}=-4t^{2}U+(4t^{2}-U^{2}+\Delta v^{2})E_{N,i}+2U\,E_{N,i}^{2}$ (S3)
In particular, the $N=2$-electron ground state energy, that is the lowest solution to Eq. (S3), readsCarrascal et al. (2015, 2016)
$E_{N}(\Delta v)=\frac{4t}{3}\bigg(u-w\sin\left(\theta+\frac{\pi}{6}\right)\bigg)$ (S4)
The angle $\theta$ is defined as
$\cos(3\theta)=\frac{u}{w^{3}}\Big(9\big(\nu^{2}-1/2\big)-u^{2}\Big)$ (S5)
with $u=U/2t$, $w=\sqrt{3(1+\nu^{2})+u^{2}}$ and $\nu=\Delta v/2t$.
Ensemble density and Fukui functions from the external potential – The ensemble is comprised of the 1-to-3 electron(s) ground-states, which all separately satisfy the variational principle
$E_{N\pm p}[v]=\underset{n_{N}\pm p}{\rm min}\Big\{F[n_{N\pm p}]+\int d{\bf r}\;v({\bf r})\;n_{N\pm p}({\bf r})\Big\}$ (S6)
with $F[n_{N\pm p}]$ the universal Hohenberg-Kohn functional. By the means of a Legendre-Fenchel transform, we can obtain an expression for the latterDeur et al. (2018):
$F[n_{N\pm p}]=\underset{v}{\rm sup}\Big\{E_{N\pm p}[v]-\int d{\bf r}\;v({\bf r})\;n_{N\pm p}({\bf r})\Big\}$ (S7)
from which we can deduce the link between density and energy-derivative
$n_{N\pm p}({\bf r})=\frac{\delta E_{N\pm p}[v]}{\delta v({\bf r})}$ (S8)
In Hubbard dimer, the Site Occupation Functional Theory (SOFT) equivalent is
$F_{N\pm p}(\Delta n_{N\pm p})=\underset{\Delta v}{\rm sup}\Big\{E_{N\pm p}[\Delta v]-\frac{\Delta v}{2}\;\Delta n_{N\pm p}\Big\}$ (S9)
with $\Delta n_{N\pm p}$ the density difference between site 1 and site 0. As the sum of density on both sites always equals $N\pm p$, we can use the density on site 0, noted $n_{N\pm p}$, as the basic variable
$\frac{\Delta n_{N\pm p}}{2}=\frac{N\pm p}{2}-n_{N\pm p}=\frac{\partial E_{N\pm p}[\Delta v]}{\partial\Delta v}$ (S10)
To obtain the ground-state energy of the 1-electron system we note it lacks electronic correlation entirely, and is thus given by the lowest orbital energy in the non-interacting 2-electron systemSenjean and Fromager (2018) with external potential $\Delta v$
$E_{N_{-}}(\Delta v)=-\sqrt{t^{2}+(\Delta v/2)^{2}}$ (S11)
while hole-particle symmetrySenjean and Fromager (2018) allows to obtain the 3-electrons ground-state energy as
$E_{N_{+}}(\Delta v)=U-\sqrt{t^{2}+(\Delta v/2)^{2}}$ (S12)
Inserting Eqs. (S4,S11,S12) in Eq. (S10) yields
$\left\{\begin{split}\;&n_{N_{-}}(\Delta v)=\frac{1}{2}+\frac{1}{4}\frac{\Delta v}{\sqrt{t^{2}+(\Delta v/2})^{2}}\\ &n_{N}(\nu)=1+\frac{2\nu}{w}\sin\left(\theta+\frac{\pi}{6}\right)+\Theta\cos\left(\theta+\frac{\pi}{6}\right)\\ &n_{N_{+}}(\Delta v)=\frac{3}{2}+\frac{1}{4}\frac{\Delta v}{\sqrt{t^{2}+(\Delta v/2})^{2}}\end{split}\right$. (S13)